A student releases a marble from the top of a ramp. The marble increases speed while on he ramp then continues across the floor. The marble travels of 140cm in 4.80s.


1
Expert's answer
2021-01-28T17:09:29-0500
"v=\frac{s}{t}\\v=\frac{1.4}{4.8}=0.292\frac{m}{s}"